# Can Carbon-Aware Electric Load Shifting Reduce Emissions? An Equilibrium-Based Analysis

### v1

@@ III-B Analysis for Different Carbon Costs
 [tons/MWh]
$u_{d}^{\intercal}P_{D}$ [\$] $\lambda\sum\limits_{d\in\mathcal{D}}c_{D,d}\cdot P_{D,d}$ [\$]
MCFD 8550 63748 3001.8 0.351 457039 0
MCDF 8550 63748 3001.8 0.351 457039 0
$[10,40]$ 8550 63748 3001.8 0.351 457039 80271.3
$[30,60]$ 8515.8 63358 2969 0.349 456301.4 138965.4
$[50,80]$ 8368.4 61677.3 2827.4 0.338 452348 189575.5
$[80,100]$ 8159.2 59986.7 2680.8 0.329 445700.2 244915.1
@@
We first observe that the two carbon-agnostic market clearing models (MCFD and MCDF) lead to the same solutions. This indicates that the consumer utility values $r_{D}$ are high enough to cause loads to consume at their highest level in the formulation with load flexibility, but no carbon costs. Next, we consider the impact of low carbon costs in the range $[10,40]$ (the third row of Table III). For these carbon costs, we observe a negligible impact on total power generation, indicating that the loads are still consuming the same amount of electricity despite occurring a carbon cost. We conclude that when carbon costs are low, they do not impact consumers’ electricity consumption. Further, we consider the case with higher carbon costs in the ranges $[30,60]$, $[50,80]$, or $[80,100]$ (lower three rows of Table III). As the carbon costs increase to these levels, the total demand decreases, and the total generation decreases accordingly. This happens because some loads choose to consume less, as the carbon-dependent utility is too low (because the carbon cost becomes too high). As the total generation decreases, the total generation cost, total carbon emissions, and average carbon emissions also decrease.
III-C Comparing against the Sequential Method
@@
We next seek to compare our results to those achieved if we use the sequential method. We run the sequential method with the same input data as we provided to the equilibrium model in Section III-B. The results for both methods under different carbon cost ranges are shown in Fig. 4.
(a) Total Generation (MWh).
(b) Total Generation Cost (\$).
(c) Total Carbon (tons).
(d) Average Carbon (tons/MWh).
(e) Total Carbon Cost (\$).
Figure 4: Comparison of equilibrium model with sequential load shifting model.
@@
We first focus on the case with carbon costs in $[10,40]$. Due to carbon costs being not high enough, both the equilibrium method and the sequential method share the same optimal solutions compared to carbon-agnostic results. With higher carbon costs in the ranges at or above $[30,60]$, the total generation is reduced due to demand reductions. We observe that the results from the equilibrium method and the sequential method are the same for carbon costs of $[30,60]$ and $[80,100]$, but different for the case with $[50,80]$. Specifically, the sequential method has lower total generation (indicating larger demand reduction) and lower average carbon emissions as compared with the equilibrium method. We next explain these differences by comparing the average carbon emissions observed by the sequential method when performing load shifting in step $2$ (which are based on the initial market clearing) with the average carbon emissions after the market has cleared again in step $3$, as well as the average carbon emissions from the equilibrium method.
@@
For the [50,80] case, consumers in the sequential model observe an average carbon emission of $\lambda_{b}=0.351$ after the initial market clearing (Step $1$). This $\lambda_{b}$ is high enough for some loads to reduce their consumption from $P_{D,d}^{\max}$ to $P_{D,d}^{\min}$. This leads to a reduction in load which lowers the average carbon emission to $\lambda_{a}=0.336$ after the market clears again (Step $3$). This $\lambda_{a}$ is low enough for some loads to want to increase their consumption again; however, they do not know this at the time of shifting and are unable to act after the market has cleared the second time (i.e., consumers conduct load shifting in Step 2 only using last clearing results but without knowing the impacts of their behaviors on the next clearing results). With the equilibrium model, the loads can “see” the impact of their demand reductions when the market is cleared. In this case, the market participants “agree” to a dispatch where the average carbon emissions $\lambda=0.338$. With this average carbon emissions, more loads choose to consume $P_{D,d}^{\max}$ as compared with the sequential load shifting case. This leads to a smaller load reduction, which corresponds to higher total emissions. However, the solution to the equilibrium model is more aligned with the objective of maximizing the (carbon-dependent) utility of the loads.

### v2

@@ 4.2. Analysis for Different Carbon Costs
 [tCO2/MWh]
MCFD 8550 63748 3001.8 0.351
MCDF 8550 63748 3001.8 0.351
$[10,40]$ 8550 63748 3001.8 0.351
$[30,60]$ 8515.8 63358 2969 0.349
$[50,80]$ 8368.4 61677.3 2827.4 0.338
$[80,100]$ 8159.2 59986.7 2680.8 0.329
@@
We first observe that the two carbon-agnostic market clearing models (MCFD and MCDF) lead to the same solutions. This indicates that the consumer utility values $r_{D}$ are high enough to cause loads to consume at their highest level in the formulation with load flexibility, but no carbon costs. Next, we consider the impact of low carbon costs in the range $[10,40]$ (the third row of Table 3). For these carbon costs, we observe a negligible impact on total power generation, indicating that the loads are still consuming the same amount of electricity despite occurring a carbon cost. We conclude that when carbon costs are low, they do not impact consumers’ electricity consumption. Further, we consider the case with higher carbon costs in the ranges $[30,60]$, $[50,80]$, or $[80,100]$ (lower three rows of Table 3). As the carbon costs increase to these levels, the total demand decreases, and the total generation decreases accordingly. This happens because some loads choose to consume less, as the carbon-dependent utility is too low (because the carbon cost becomes too high). As the total generation decreases, the total generation cost, total carbon emissions, and average carbon emissions also decrease.
4.3. The Impacts of Carbon Costs on Electricity Prices and Generation Profits
The introduction of carbon-dependent consumer utility will impact market clearing results, leading to different electricity prices faced by both generators and consumers as well as different generation profits. We next analyze how carbon-dependent utility, represented through carbon costs, impacts electricity prices and generation profits compared to those obtained from a standard market clearing (i.e., where all carbon costs are zero). To assess this, we assign consumer carbon costs $c_{D}$ in a range of $[30,60]$ \$/MWh, and then randomly assign zero carbon costs to 25% of the consumers to simulate carbon-agnostic consumers. For consumers’ demand flexibility, we use 120% of values in the RTS-GMLC system as the maximums.
In this scenario, the introduction of carbon costs causes consumers to use less electricity, and the total electricity generation decreases from 10,260 MWh in the standard benchmark to 10,093.6 MWh (-1.6%). The carbon emissions fall from 3,900.6 tCO2 to 3,785.5 tCO2 (-2.9%), while the overall generation cost drops from \$77,910.4 to \$75,187.8 (-3.6%). These results suggest that a small reduction in loads leads to a larger reduction in emissions and cost.
To show how the load shifting impacts prices and profits of individual consumers and generators, Fig. 2(d) plots the profit for generators with the carbon-aware and standard market clearing (left) and the difference in LMPs for different consumers based on results from our equilibrium model and standard model (right). In addition, Fig. 2 illustrates the geographical variation in LMPs at nodes throughout the system and shows where the congested lines are located.
From Fig. 2(d).a, we observe that most generators experience a reduction in their profits as consumers reduce their demand. This decline in generator profits is more pronounced for carbon-intensive generators than for low-carbon generators. Further, Fig 2(d).b illustrates that the reduction in demand from carbon-sensitive consumers leads to lower LMPs for most consumers. This price reduction is primarily driven by the alleviation of transmission congestion, as highlighted in Fig. 2. Although the same set of lines (27, 85, 109, and 119) remain congested in both carbon-aware and carbon-agnostic scenarios, congestion levels are lower under carbon-aware conditions. This is reflected in the dual variables of constraint (3b), which are -43.9, 2.2, 7.9, and 1.3 in the carbon-aware case, compared to -51.4, 2.6, 9.3, and 1.5 in the carbon-agnostic case. Furthermore, Fig. 2 demonstrates that consumers with higher LMPs (shown by dark purple triangles) experience greater reductions in LMPs (shown by light blue circles) under the carbon-aware scenario than those with lower LMPs (light purple triangles are mostly linked to dark blue circles).
(d) Impact of carbon costs on generation profits and electricity prices.
Figure 2. Changes in LMPs with the introduction of carbon cost. Each circle represents a bus node, with the color intensity representing the difference between carbon-aware and standard LMPs. The squares represent generators, with the color intensity reflecting their carbon emission intensity factors. The triangles attached to each node represent consumers, with the color intensity representing their respective carbon costs. Congested transmission lines are highlighted by the red color (the uncongested lines are in gray).
4.4. Comparing against the Sequential Method
@@
We next seek to compare our results to those achieved if we use the sequential method. We run the sequential method with the same input data as we provided to the equilibrium model in Section 4.2. The results for both methods under different carbon cost ranges are shown in Fig. 2(e).
(e) Comparison of the equilibrium model with the sequential load shifting model.
@@
We first focus on the case with carbon costs in $[10,40]$. Due to carbon costs not being high enough, both the equilibrium method and the sequential method share the same optimal solutions compared to carbon-agnostic results. With higher carbon costs in the ranges at or above $[30,60]$, the total generation is reduced due to demand reductions. We observe that the results from the equilibrium method and the sequential method are the same for carbon costs of $[30,60]$ and $[80,100]$, but different for the case with $[50,80]$. Specifically, the sequential method has lower total generation (indicating larger demand reduction) and lower average carbon emissions as compared with the equilibrium method. We next explain these differences by comparing the average carbon emissions observed by the sequential method when performing load shifting in step $2$ (which is based on the initial market clearing) with the average carbon emissions after the market has cleared again in step $3$, as well as the average carbon emissions from the equilibrium method.
@@
For the [50,80] case, consumers in the sequential model observe an average carbon emission of $\lambda_{b}=0.351$ after the initial market clearing (Step $1$). This $\lambda_{b}$ is high enough for some loads to reduce their consumption from $P_{D,d}^{\max}$ to $P_{D,d}^{\min}$. This leads to a reduction in load, which lowers the average carbon emission to $\lambda_{a}=0.336$ after the market clears again (Step $3$). This $\lambda_{a}$ is low enough for some loads to want to increase their consumption again; however, they do not know this at the time of shifting and are unable to act after the market has cleared the second time (i.e., consumers conduct load shifting in Step 2 only using last clearing results but without knowing the impacts of their behaviors on the next clearing results). With the equilibrium model, the loads can “see” the impact of their demand reductions when the market is cleared. In this case, the market participants “agree” to a dispatch where the average carbon emissions $\lambda=0.338$. With these average carbon emissions, more loads choose to consume $P_{D,d}^{\max}$ as compared with the sequential load shifting case. This leads to a smaller load reduction, which corresponds to higher total emissions. However, the solution to the equilibrium model is more aligned with the objective of maximizing the (carbon-dependent) utility of the loads.